OpenStudy (anonymous):

Indicate the equation of the given line in standard form. The line through the midpoint of and perpendicular to the segment joining points (1, 0) and (5, -2).

OpenStudy (sgadi):

slope of given line sigment = -2/4=-0.5 solve of the required line is =-1/(-0.5)=2 mid point of line sigment = (3,-1) the required line is having the equation of line with slope 2 and passing though point (3,-1) y=2x-7